OpenStudy (anonymous):

what is 2 and 2/3 - 3/5 a.7/15 b.1 2/5 c.2 1/15 d.4 4/15

OpenStudy (anonymous):

\[\huge 2\frac{ 2 }{ 3 }-\frac{ 3 }{ 5 }~?\]is this your question?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

first change the mixed fraction in to fraction \[2\frac{ 2 }{ 3 } = \frac{ 8 }{ 3 }\] then find the LCD for 3 and 5 their LCD would be 15 \[\color{red}{5}*\frac{ 8 }{ 3 } - \color{red}{3}*\frac{ 3 }{ 5 }\] \[\frac{ 40 }{ 15 }-\frac{9}{15} = \frac{31}{15}\]

OpenStudy (anonymous):

then turn \[\frac{ 31 }{ 15 }\] into mixed fraction

OpenStudy (anonymous):

which would be \[2\frac{ 1 }{ 15 }\]

OpenStudy (anonymous):

thank you very much i suck at fractions

OpenStudy (anonymous):

Anytime :)